# Leveraging Discrete Choice Experiments for User-Centric Requirements Prioritization in mHealth Applications

### v1

@@ 5.2.3 How do user preferences differ among various user demographics
Understanding how user preferences differ across demographics, including age, health conditions, and mechanisms for managing stress related to diseases, is critical to designing inclusive and effective mHealth solutions. Traditionally, mHealth applications have primarily focused on younger and healthier audiences [boulos2011smartphones], often neglecting the needs of older adults or individuals with chronic diseases. This lack of inclusive design risks exacerbating health disparities and undermining the goal of universal access to healthcare services [van2022digital]. Insights from DCE can also guide inclusive design practices, ensuring that mHealth apps are accessible and effective for all users. By focusing on the requirements of underrepresented groups, including elderly individuals or those with serious chronic disease, DCEs can work towards diminishing health inequalities and foster fair access to medical services. Beyond mHealth, DCEs have broader applicability in domains such as education, finance, and e-commerce, where understanding user preferences is critical to designing effective systems. For example, in education, DCEs can help prioritize features for personalized learning platforms [marienko2020personalization], while in e-Commerce, they can inform the design of user-friendly interfaces.
@@ 6 Threats to validity
Internal Validity. First, the study relied on self-reported data, which can introduce biases such as social desirability or inaccurate recall [chan2010so]. Though there were efforts to mitigate biases, these may have influenced participants’ stated preferences, and as with all stated preference methods, the results reflect hypothetical choices in a controlled survey context, which may not fully align with their revealed preferences in real-world use. Furthermore, variations in the way the participants interpreted the attributes may have occurred, despite a comprehension check carried out before the DCE. We also included a random tenth question that mirrors the previous to assess whether responses showed consistency by choosing the same adaptation design. Furthermore, while the CIs in Table 2 provide insight into participant understanding, this variation constrains the strength of our conclusions. An excessive number of choice sets or overly complex attribute combinations in a DCE can result in participant fatigue or cognitive overload, potentially causing random or inconsistent responses. To address this, we limited the survey to a selection of choice sets, implementing a fractional factorial design [soekhai2019discrete, bridges2011conjoint]. The sequence in which choice sets or attributes are presented could affect responses [bridges2011conjoint], and participants could alter their decision-making approach as the experiment progresses, leading to inconsistent responses. To mitigate order effects, we randomized both the order of attributes and the sequence of questions each participant received upon survey entry.
@@
External Validity. The sample size, while sufficient for statistical analysis, may not fully represent the diversity of people with chronic diseases, particularly in terms of cultural, socioeconomic, and technological backgrounds. Furthermore, 62% of participants had cardiometabolic and endocrine conditions (Table 2), which may limit the generalizability of findings to populations with other health conditions. Similarly, the participants’ familiarity with technology may have influenced their preferences, restricting broader applicability to less tech-savvy groups. Utilizing Prolific as a recruitment platform could potentially lead to sampling bias. Despite Prolific’s reputation for a varied set of participants, its users generally possess technological proficiency and a familiarity with online surveys. This might skew the sample towards individuals who already found mHealth applications easy to use. Although Prolific implements quality control measures, the platform relies on self-reported health information, which may introduce inaccuracies or misclassification in participant eligibility. Furthermore, integrating Prolific’s responses with those obtained through alternative recruitment methods could result in heterogeneity of the sample, as participants sourced from various channels could vary in their demographic or psychographic characteristics.
@@
Construct Validity. The DCE methodology, while effective in understanding trade-offs and preferences, simplifies real-world decision-making by isolating attributes and limiting context [de2012discrete, soekhai2019discrete]. The simplification and hypothetical nature of DCEs may not adequately reflect the intricate requirements of users in adaptive mHealth applications, raising doubts about whether the choices in DCEs align with real-world decisions, known as the intention-behavior gap [ajzen1991theory]. Although participants may assume that they would choose a scenario from a set of options, in real-world contexts, additional factors like privacy concerns regarding adaptation or user support of adaptation [trewin2000configuration, hamidi2018should] could influence their decisions.
@@
Conclusion Validity. Our sample size was sufficient for the primary analyses, but it limited the ability to explore interactions between demographic factors (e.g., age and health condition) without substantially reducing statistical power. Further stratification would have produced smaller subgroups, increasing the risk of unstable estimates and Type II errors. Future studies with larger and more diverse samples are needed to examine these interactions with greater reliability.
@@ 7 Conclusion
This study explored user preferences for Adaptive User Interfaces (AUIs) in mHealth applications, focusing on individuals with chronic diseases through a Discrete Choice Experiment (DCE). By prioritizing attributes such as usability, controllability over adaptations, frequency of adaptations, and granularity of adaptations, the findings underscore the importance of designing mHealth applications that align with user needs and preferences. This study contributes to the advancement of adaptation design for mHealth applications by offering a robust framework for understanding and addressing competing user requirements. Through a data-driven approach to requirements prioritization, we quantify user preferences and assess the relative importance of key attributes. By analyzing preference heterogeneity, we reveal how preferences vary across demographic and behavioral subgroups, highlighting the trade-offs users are willing to make. Our findings offer practical insights for designing next-generation adaptive mHealth applications and set the stage for further research in software engineering, particularly in user-centered adaptation strategies and requirements prioritization.
@@ Acknowledgements
Wang and Grundy are supported by ARC Laureate Fellowship FL190100035. Our recruitment was partly supported by Join Us, a National, not-for-profit and disease agnostic research register founded by UNSW and The George Institute. We express our gratitude to all participants who completed our survey and those who provided valuable feedback to help refine it. We would like to express our sincere gratitude to Dawn Mischewski for her invaluable assistance in refining the survey design.
@@
During the writing process of this work the authors used ChatGPT in order to improve and rephrase written paragraphs. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ 5.2.3 How do user preferences differ among various user demographics
Understanding how user preferences differ across demographics, including age, health conditions, and mechanisms for managing stress related to diseases, is critical to designing inclusive and effective mHealth solutions. Traditionally, mHealth applications have focused mainly on younger and healthier audiences [13], often neglecting the needs of older adults or individuals with chronic diseases. This lack of inclusive design risks exacerbating health disparities and undermining the goal of universal access to healthcare services [92]. Insights from DCE can also guide inclusive design practices, ensuring that mHealth apps are accessible and effective for all users. By focusing on the requirements of underrepresented groups, including elderly individuals or those with serious chronic disease, DCEs can work towards diminishing health inequalities and foster fair access to medical services. Beyond mHealth, DCEs have broader applicability in domains such as education, finance, and e-commerce, where understanding user preferences is critical to designing effective systems. For example, in education, DCEs can help prioritize features for personalized learning platforms [64], while in e-Commerce, they can inform the design of user-friendly interfaces.
@@ 6 Threats to validity
Internal Validity. First, the study relied on self-reported data, which can introduce biases such as social desirability or inaccurate recall [21]. Although there were efforts to mitigate biases, these may have influenced the’ stated preferences of the participants, and as with all stated preference methods, the results reflect hypothetical choices in a controlled survey context, which may not fully align with their revealed preferences in real-world use. Furthermore, variations in the way the participants interpreted the attributes may have occurred, despite a comprehension check carried out before the DCE. We also included a random tenth question that mirrors the previous to assess whether responses showed consistency by choosing the same adaptation design. Furthermore, while the CIs in Table 2 provide insight into participant understanding, this variation constrains the strength of our conclusions. An excessive number of choice sets or overly complex attribute combinations in a DCE can result in participant fatigue or cognitive overload, potentially causing random or inconsistent responses. To address this, we limited the survey to a selection of choice sets, implementing a fractional factorial design [86, 16]. The sequence in which choice sets or attributes are presented could affect responses [16], and participants could alter their decision-making approach as the experiment progresses, leading to inconsistent responses. To mitigate order effects, we randomly assigned both the order of attributes and the sequence of questions each participant received upon survey entry.
@@
External Validity. The sample size, while sufficient for statistical analysis, may not fully represent the diversity of people with chronic diseases, particularly in terms of cultural, socioeconomic, and technological backgrounds. Furthermore, 62% of the participants had cardiometabolic and endocrine conditions (Table 2), which may limit the generalizability of the findings to populations with other health conditions. Similarly, the participants’ familiarity with technology may have influenced their preferences, restricting broader applicability to less tech-savvy groups. Using Prolific as a recruitment platform could potentially lead to sampling bias. Despite Prolific’s reputation for a varied set of participants, its users generally possess technological proficiency and a familiarity with online surveys. This might skew the sample towards people who already found mHealth applications easy to use. Although Prolific implements quality control measures, the platform relies on self-reported health information, which can introduce inaccuracies or misclassification in participant eligibility. Furthermore, integrating Prolific’s responses with those obtained through alternative recruitment methods could result in heterogeneity of the sample, as participants sourced from various channels could vary in their demographic or psychographic characteristics.
@@
Construct Validity. The DCE methodology, while effective in understanding trade-offs and preferences, simplifies real-world decision-making by isolating attributes and limiting context [25, 86]. The simplification and hypothetical nature of DCEs may not adequately reflect the intricate requirements of users in adaptive mHealth applications, raising doubts about whether the choices in DCEs align with real-world decisions, known as the intention-behavior gap [5]. Although participants may assume that they would choose a scenario from a set of options, in real-world contexts, additional factors, such as privacy concerns regarding adaptation or user support of adaptation [90, 39], could influence their decisions.
@@
Conclusion Validity. Our sample size was sufficient for the primary analyzes, but it limited the ability to explore interactions between demographic factors (e.g., age and health condition) without substantially reducing statistical power. Further stratification would have produced smaller subgroups, increasing the risk of unstable estimates and Type II errors. Future studies with larger and more diverse samples are needed to examine these interactions with greater reliability.
@@ 7 Conclusion
This study contributes a data-driven approach to modeling user preferences and trade-offs in adaptive mHealth interface design, using a Discrete Choice Experiment (DCE) with individuals managing chronic diseases. By prioritizing attributes such as usability, controllability over adaptations, frequency of adaptations, and granularity of adaptations, the findings underscore the importance of designing mHealth applications that align with user needs and preferences. This study contributes to the advancement of adaptation design for mHealth applications by offering a robust framework for understanding and addressing competing user requirements. Through a data-driven approach to requirements prioritization, we quantify user preferences and assess the relative importance of key attributes. By analyzing preference heterogeneity, we reveal how preferences vary across demographic and behavioral subgroups, highlighting the trade-offs users are willing to make. These insights inform the design of more inclusive and responsive mHealth applications. Future work could extend this framework to support dynamic personalization strategies and integrate user-driven prioritization into broader requirements engineering processes for adaptive health technologies.
@@ Acknowledgements
Wang and Grundy are supported by the ARC Laureate Fellowship FL190100035. Our recruitment was partly supported by Join Us, a National, not-for-profit and disease agnostic research register founded by UNSW and The George Institute. We express our gratitude to all participants who completed our survey and those who provided valuable feedback to help refine it. We express our sincere gratitude to Dawn Mischewski for her invaluable assistance in refining the survey design.
@@
During the writing process of this work, the authors used ChatGPT to improve and rephrase written paragraphs. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.